# TripNet: Learning Large-scale High-fidelity 3D Car Aerodynamics with Triplane Networks

### v1

@@ Abstract
Computational Fluid Dynamics (CFD) simulations are essential in product design, providing insights into fluid behavior around complex geometries in aerospace and automotive applications. However, high-fidelity CFD simulations are computationally expensive, making rapid design iterations challenging. To address this, we propose TripNet, Triplane CFD Network, a machine learning-based framework leveraging triplane representations to predict the outcomes of large-scale, high-fidelity CFD simulations with significantly reduced computation cost. Our method encodes 3D geometry into compact yet information-rich triplane features, maintaining full geometry fidelity and enabling accurate aerodynamic predictions. Unlike graph- and point cloud-based models, which are inherently discrete and provide solutions only at the mesh nodes, TripNet allows the solution to be queried at any point in the 3D space. Validated on high-fidelity DrivAerNet and DrivAerNet++ car aerodynamics datasets, TripNet achieves state-of-the-art performance in drag coefficient prediction, surface field estimation, and full 3D flow field simulations of industry-standard car designs. By utilizing a shared triplane backbone across multiple tasks, our approach offers a scalable, accurate, and efficient alternative to traditional CFD solvers.
Machine Learning, ICML
@@ 1 Introduction
Computational simulations have become a cornerstone of engineering and scientific research in recent decades, allowing for the investigation and study of complex physics efficiently and cost-effectively. These simulations solve the partial differential equations (PDEs) to model real-world behavior. Mesh-based simulation approaches such as the finite volume method (FVM)  (Blazek, 2015; Patankar, 2018), finite element method (FEM)  (Bathe, 2006; Belytschko et al., 2014), and finite difference method (FDM)  (Pletcher et al., 2012) have been widely adopted in fields such as aerospace and automotive engineering for computational fluid dynamics and crash simulations. These methods significantly reduce the need for expensive and time-consuming wind tunnel experiments and collision tests.
@@
Simulations are pivotal in driving advancements in science and engineering. Yet, executing large-scale, high-fidelity simulations demands considerable computational resources and time, often extending from several days to weeks per simulation. To address this issue and accelerate mesh-based simulations, machine learning has emerged as a transformative tool to accelerate traditional simulations by offering faster approximations without sacrificing accuracy. Among the promising machine learning approaches, MeshGraphNets (Pfaff et al., 2020; Nabian, 2024) have shown considerable potential. These models leverage graph networks to learn high-dimensional scientific simulation data. The model undergoes training with simulation data through several rounds of message-passing along mesh-edges and world-edges, enabling it to accurately predict dynamics across aerodynamics, structural mechanics, and cloth simulations. However, the scalability of MeshGraphNets is somewhat limited to a few thousand nodes due to the high computational costs associated with building graphs for high-resolution meshes. While MeshGraphNets and similar models (Li et al., 2024) have demonstrated significant progress, they face challenges in scaling to large datasets and high-resolution meshes due to computational and memory constraints. Furthermore, alternative representations such as images (Song et al., 2023; Elrefaie et al., 2024c), point clouds (Qi et al., 2017a, b), and parametric models (Elrefaie et al., 2024b) often suffer from either a loss of fine geometric details or excessive computational overhead.
To overcome these challenges, we propose a 3D geometric deep learning model that leverages the triplane representation for tasks such as aerodynamic drag, surface pressure, and 3D flow field prediction. The triplane representation provides an implicit and continuous representation of the 3D geometry, allowing the model to query any point within the flow field domain, unlike traditional discrete methods, which require computations over the entire representations such as graphs and point cloud, triplanes enable efficient point-wise querying within the domain. The triplanes are processed by a convolutional backbone, allowing the model to learn how global geometry influences local aerodynamic behavior. This structured representation leverages pure convolutional operations for global drag predictions and a hybrid convolutional-MLP architecture for field prediction in a query-based fashion. By capturing rich geometric features while remaining compact and efficient, triplanes offer a powerful foundation for high-fidelity surrogate modeling. Moreover, since triplanes are extensively employed by generative model (Shue et al., 2023; Sun et al., 2023; Zou et al., 2024), the surrogate models we developed can be seamlessly integrated into various generative AI frameworks for performance-aware 3D generation.
@@
    TripNets enable aerodynamic predictions for industry-standard cars with remarkable speed and accuracy. Using a single Nvidia RTX 4090 GPU, the method accelerates CFD simulations by several orders of magnitude, reducing the simulation time from 10 hours to seconds. Key performance highlights include drag coefficient tasks less than 0.002 seconds, surface field prediction for 0.5 million nodes under 0.1 seconds, and full 3D velocity filed prediction (12 million cells) in 2 seconds compared to tradition traditional CFD simulation. [^1] For a steady-state RANS simulation on an HPC cluster with 256 CPU cores for 7,000 iterations.
@@
These contributions highlight the effectiveness and scalability of our approach for high-fidelity aerodynamic analysis. To the best of the authors’ knowledge, this is the first implementation of the triplane representation for regression tasks, as all existing methods have exclusively utilized it for generative modeling.
(a) Fitting triplane features with pre-trained MLP to reconstruct the occupancy field.
(b) Applications of TripNets on different CFD tasks.
(c) Hybrid model combining U-Net and MLP for surface and volume field prediction.
@@
Figure 1: Illustration of Triplane Networks and their role in CFD tasks. The figure demonstrates how triplane features encode spatial information for aerodynamic analysis, showcasing their applications in predicting drag coefficients, surface pressure, wall shear stress, and full 3D flow fields. Additionally, it highlights the integration of U-Net and MLP architectures for efficient surface and volumetric flow field predictions.
2 Related Work
RANS for Initial Design Stages
Reynolds-Averaged Navier-Stokes (RANS) simulations are a fundamental approach in industrial fluid dynamics, particularly valuable in the early and conceptual design stages due to their balance of accuracy and computational efficiency. Recent datasets for aerodynamic analysis have employed RANS simulations, showcasing a wide range of dataset sizes and mesh resolutions. These include 500 designs in Usama et al. (Usama et al., 2021), 889 in Umetani et al. (Umetani & Bickel, 2018), 1,000 in Gunpinar et al. (Gunpinar et al., 2019), and 1,400 in Remelli et al. (Remelli et al., 2020). Larger datasets include 2,474 designs in Song et al. (Song et al., 2023) and 1,121 in Trinh et al. (Trinh et al., 2024), with mesh resolutions ranging from 130k in Gunpinar et al. (Gunpinar et al., 2019) to 3M in Song et al. (Song et al., 2023). The most recent DrivAerNet (Elrefaie et al., 2024a) and DrivAerNet++ (Elrefaie et al., 2024d) datasets were published to address the trade-off between large dataset sizes required for training deep learning models and the simulation fidelity necessary for industrial applications. These datasets comprise 8,000 designs with high-resolution meshes, providing an unparalleled level of detail for evaluating machine learning-based CFD approaches. This combination ensures robust performance benchmarking and facilitates the development of more accurate predictive models.
Deep Learning Models for Simulation:
Jacob et al. (Jacob et al., 2021) used Signed Distance Fields (SDF) to represent the input geometries and a modified U-Net to predict both the drag coefficient and velocity fields simultaneously. Neural operator is a family of machine learning models designed the approximate the solutions of PDEs. Unlike neural networks that map inputs to outputs, neural operators are trained to approximate the mapping between function spaces. Among the popular architectures, the Graph Neural Operator (GNO)  (Li et al., 2024) adopts a graph representation for kernel integration, making it highly suitable for handling complex geometries and irregular domains. The Fourier Neural Operator (FNO)  (Li et al., 2020) utilizes the Fourier transform to represent the kernel integration and solves PDEs in the spectral domain, achieving computational efficiency on regular grids by leveraging Fast Fourier Transforms (FFT). Transolver: A fast transformer-based solver for partial differential equations (PDEs) on general geometries (Wu et al., 2024a) operates on mesh points $N$, ranging from $10242^{10}=$ 1024 to $7682^{15}=32,$ 768. While it demonstrated superior performance compared to linear transformer models, scaling to significantly larger datasets—such as the DrivAerNet++ dataset, which is three orders of magnitude larger—remains a challenge.
Different representations on 3D shapes have been used by machine learning models for CFD simulation, particularly the drag coefficient prediction. Song et al (Song et al., 2023) adopts a new 3D representation using stacked depth and normal renderings. Six single views, front, rear, top, bottom, right and left, are generated and integrated into a single image to capture the geometric features comprehensively. CNN-based and transformer-based models are used to learn the features from the 2D representation and predict drag coefficients.
Graph Neural Networks (GNNs) have emerged as one of the most popular approaches for learning-based CFD simulations due to their ability to handle irregular geometries and unstructured meshes. However, several challenges remain in their application to large-scale, high-fidelity aerodynamic datasets. MeshGraphNets (MGNs) have been extensively studied for their generalization capabilities in fluid dynamics, particularly for unseen geometries (Schmöcker et al., 2024). A comparative analysis of computational performance on both CPU and GPU systems demonstrated a 6.5x speedup in inference on a CPU and a 100.8x speedup on a GPU, significantly reducing training and inference times for replicating full CFD simulations. However, this study primarily focuses on fluid simulations for standard geometries, utilizing meshes with $\mathcal{O}(10^{3})$ nodes. Two key limitations of this approach are (1) the use of relatively simple geometries (e.g., cylinders) and (2) the restricted number of nodes, again at $\mathcal{O}(10^{3})$. To enhance GNN-based methods for CFD simulations, an alternative approach introduces Shortest Vector (SV) and Directional Integrated Distance (DID) as global geometry representations, along with Finite Volume Features (FVF) as node and edge attributes (Jessica et al., 2023). These innovations allow graph convolutions to more closely model the finite volume simulation method while eliminating the need for message-passing to collect global geometry information. Additionally, residual training with low-resolution data is leveraged to improve flow field prediction accuracy, outperforming state-of-the-art methods such as MeshGraphNet and Graph U-Net, reducing predictive error by up to 41%.
Implicit and Neural Representations:
Implicit and neural representations (Serrano et al., 2023; Catalani et al., 2024) offer a continuous alternative to traditional discrete approaches in CFD simulations. Unlike GNNs and point cloud-based models, which provide solutions only at discrete mesh nodes, implicit representations enable querying the solution at any point in space (and time, if included as an input feature). This flexibility allows for higher-resolution predictions without being constrained by the mesh structure, making them particularly advantageous for modeling complex aerodynamic flows. de Vito et al. (de Vito et al., 2024) presents an implicit neural representation framework for accurate 3D CFD flow field prediction, leveraging a coordinate-based MLP (backbone-net) to provide a discretization-agnostic, infinite-resolution representation of flow fields for domains with arbitrary topology. Combined with a hyper-net that maps blade surface geometry to the backbone-net’s parameters, this approach offers a memory-efficient, highly accurate proxy to CFD solvers, demonstrating strong potential for industrial applications like wing and turbine blade design.
@@ 3.1 Triplane Representation
Triplane representation has recently emerged as a powerful approach in 3D representation learning for scene reconstruction (Wu et al., 2024b), 3D geometry (Zhang et al., 2024) and avatar generation  (Wang et al., 2023; Sun et al., 2023; Zhang et al., 2025) because of its balance between computational efficiency and ability to capture fine-grained details in 3D objects. Three axis-aligned and orthogonal feature planes are used to encode 3D information, offering a compact but expressive way to handle complex 3D shapes and textures. Models such as GANs and diffusion models can be trained on these triplane representations to generate 3D neural fields with high diversity and quality (Shue et al., 2023; Chan et al., 2022). Building on this foundation, the Roll-out diffusion network (Rodin) (Wang et al., 2023) rolls out the triplanes into a single 2D feature plane and performs 3D-aware diffusion to generate 3D digital avatars. Compared to 1D latent vector, voxel, or grid-based representations, triplane representations offer greater memory efficiency while preserving spatial information, providing a 3D-aware encoding that better captures geometric and structural details.
In this work, we aim to convert the 3D car designs to triplanes for TripNet training. Firstly we build the occupancy field for each 3D object by sampling 5 million points uniformly in the domain and another 5 million points near the surface. Next, we followed the two-stage fitting method NFD (Shue et al., 2023) used to get the triplanes. Each 3D object is represented by triplanes $\mathbf{P}_{xy},\mathbf{P}_{yz},\mathbf{P}_{xz}\in\mathbb{R}^{C\times H\times W}$. In the fitting, we project the coordinate $\mathbf{x}\in\mathbb{R}^{3}$ onto the triplanes and aggregate the resulting feature vectors to predict the value of occupancy field through an MLP decoder ($MLP_{\Phi}$ with parameters $\Phi$). The occupancy function $occ:R^{3}\rightarrow\{0,1\}$ is defined as:
@@
To mitigate potential compression artifacts in triplane-based 3D representations, we draw inspiration from Neural Field Diffusion (NFD) (Shue et al., 2023), where regularization terms are introduced to simplify the data manifold that the diffusion model must learn. It has been shown that these regularization terms effectively reduce significant artifacts in 3D asset generation. We incorporate the same regularization terms in our method to enhance the quality of the triplanes. For further details, interested are encouraged to refer to the original paper (Shue et al., 2023). The training objective of the 3D reconstruction task for triplane fitting is as follows:
@@ 3.2 Drag Coefficient Task
To predict the drag coefficient ($C_{d}$), we employ a lightweight convolutional neural network (CNN) as shown in Figure. 1(a) designed to efficiently process triplane representations. The input to our model consists of raw triplane features $\mathbf{P}=[\mathbf{P}_{xy},\mathbf{P}_{yz},\mathbf{P}_{xz}]\in\mathbb{R}^{3\times 32\times 128\times 128}$. Since each feature plane has a channel depth of 32, stacking the three triplanes and taking it as the input would make the model large and inference slow. We apply channel-wise $max$, $min$, and $mean$ pooling on the raw triplane. Each pooling operator $\Theta:\mathbb{R}^{32\times 128\times 128}\rightarrow\mathbb{R}^{3\times 128\times 128}$ compresses the channel depth of original triplane $\mathbf{P}_{ij}$ from 32 to 3. Three feature planes $\mathbf{\tilde{P}}_{ij}$ where $ij\in\{xy,yz,xz\}$ are stacked channel-wise for drag coefficient prediction:
@@
Table 1: Performance comparison on drag coefficient prediction across various models using different data representations, including point clouds, graphs, and our triplane representation. The evaluation is conducted on both the DrivAerNet test set, which includes 600 industry-standard car designs, and the DrivAerNet++ test set, which includes 1200 industry-standard car designs. Models are ranked based on $R^{2}$ performance.
(×$(\times 10^{-5})$ (×$(\times 10^{-3})$ (×$(\times 10^{-2})$
DrivAerNet PointNet (Qi et al., 2017a) 12.0 8.85 10.18 0.826
GCNN (Kipf & Welling, 2016) 10.7 7.17 10.97 0.874
PointNet++ (Qi et al., 2017b) 7.813 6.755 3.463 0.896
RegDGCNN (Elrefaie et al., 2024a) 8.01 6.91 8.80 0.901
PointBERT (Yu et al., 2022) 6.334 6.204 2.767 0.915
DeepGCN (Li et al., 2019) 6.297 6.091 3.070 0.916
MeshGraphNet (Pfaff et al., 2020) 6.0 6.08 2.965 0.917
AssaNet (Qian et al., 2021) 5.433 5.81 2.39 0.927
PointNeXt (Qian et al., 2022) 4.577 5.2 2.41 0.939
FIGConvNet (Choy et al., 2025) 3.225 4.423 2.134 0.957
TripNet (Ours) 2.602 4.030 1.268 0.972
DrivAerNet++ GCNN (Kipf & Welling, 2016) 17.1 10.43 15.03 0.596
RegDGCNN (Elrefaie et al., 2024a) 14.2 9.31 12.79 0.641
PointNet (Qi et al., 2017a) 14.9 9.60 12.45 0.643
TripNet (Ours) 9.1 7.17 7.70 0.957

### v2

@@ Abstract
Surrogate modeling has emerged as a powerful tool to accelerate Computational Fluid Dynamics (CFD) simulations. Existing 3D geometric learning models based on point clouds, voxels, meshes, or graphs depend on explicit geometric representations that are memory-intensive and resolution-limited. For large-scale simulations with millions of nodes and cells, existing models require aggressive downsampling due to their dependence on mesh resolution, resulting in degraded accuracy. We present TripNet, a triplane-based neural framework that implicitly encodes 3D geometry into a compact, continuous feature map with fixed dimension. Unlike mesh-dependent approaches, TripNet scales to high-resolution simulations without increasing memory cost, and enables CFD predictions at arbitrary spatial locations in a query-based fashion, independent of mesh connectivity or predefined nodes. TripNet achieves state-of-the-art performance on the DrivAerNet and DrivAerNet++ datasets, accurately predicting drag coefficients, surface pressure, and full 3D flow fields. With a unified triplane backbone supporting multiple simulation tasks, TripNet offers a scalable, accurate, and efficient alternative to traditional CFD solvers and existing surrogate models.
@@ 1 Introduction
Computational simulations have become a cornerstone of engineering and scientific research in recent decades, allowing for the investigation and study of complex physics efficiently and cost-effectively. These simulations solve the partial differential equations (PDEs) to model real-world behavior. Mesh-based simulation approaches such as the finite volume method (FVM)  Blazek (2015); Patankar (2018), finite element method (FEM)  Bathe (2006); Belytschko et al. (2014), and finite difference method (FDM)  Pletcher et al. (2012) have been widely adopted in fields such as aerospace and automotive engineering for computational fluid dynamics and crash simulations. These methods significantly reduce the need for expensive and time-consuming wind tunnel experiments and collision tests.
@@
Simulations are pivotal in driving advancements in science and engineering. Yet, executing large-scale, high-fidelity simulations demands considerable computational resources and time, often extending from several days to weeks per simulation. To address this issue and accelerate mesh-based simulations, machine learning has emerged as a transformative tool to accelerate traditional simulations by offering faster approximations without sacrificing accuracy. Among the promising machine learning approaches, MeshGraphNets Pfaff et al. (2020); Nabian (2024) have shown considerable potential. These models leverage graph networks to learn high-dimensional scientific simulation data. The model undergoes training with simulation data through several rounds of message-passing along mesh-edges and world-edges, enabling it to accurately predict dynamics across aerodynamics, structural mechanics, and cloth simulations. However, the scalability of MeshGraphNets is somewhat limited to a few thousand nodes due to the high computational costs associated with building graphs for high-resolution meshes. While MeshGraphNets and similar models Li et al. (2024) have demonstrated significant progress, they face challenges in scaling to large datasets and high-resolution meshes due to computational and memory constraints. Furthermore, alternative representations such as images Song et al. (2023); Elrefaie et al. (2024c), point clouds Qi et al. (2017a, b), and parametric models Elrefaie et al. (2024a) often suffer from either a loss of fine geometric details or excessive computational overhead. Related work is presented in the appendix due to page limitation.
To overcome these challenges, we propose a 3D geometric deep learning model that leverages the triplane representation for tasks such as aerodynamic drag, surface pressure, and 3D flow field prediction. Triplanes offer a compact yet expressive structure that is particularly well-suited for CFD tasks. This is because aerodynamic quantities in CFD—such as pressure, wall shear stress, and velocity fields—often exhibit strong anisotropy and boundary-layer behavior aligned with surfaces. Triplane representations naturally encode such surface-oriented features across three orthogonal planes, enabling efficient learning of spatially distributed physical fields without requiring dense 3D volumes or explicit mesh connectivity. This structure strikes a balance between resolution, memory efficiency, and spatial context, making it especially advantageous for high-fidelity CFD prediction. Moreover, since triplanes are extensively employed by generative model Shue et al. (2023); Sun et al. (2023); Zou et al. (2024), the surrogate models we developed can be seamlessly integrated into various generative AI frameworks for performance-aware 3D generation.
@@
    TripNets enable aerodynamic predictions for industry-standard cars with remarkable speed and accuracy. Using a single Nvidia RTX 4090 GPU, the method accelerates CFD simulations by several orders of magnitude, reducing the simulation time from 10 hours to seconds. Key performance highlights include drag coefficient tasks less than 0.01 seconds, surface field prediction for 0.5 million nodes under 0.1 seconds, and full 3D velocity field prediction (12 million cells) in 2 seconds compared to traditional CFD simulation. [^1] For a steady-state RANS simulation on an HPC cluster with 256 CPU cores for 7,000 iterations. TripNet outperforms prior SOTA models in accuracy on a variety of simulation tasks while being more memory efficient and faster than prior SOTA models: TripNet outperforms all prior models in drag coefficient prediction, achieving an $R^{2}$ score of 0.972 on DrivAerNet and 0.957 on DrivAerNet++; For surface pressure prediction, TripNet achieves the lowest $\ell_{2}$ error of 20.25% on DrivAerNet and 20.05% on DrivAerNet++. For surface wall shear stress prediction, it also attains the lowest error of 22.07% on DrivAerNet++ and the second-lowest error of 21.80% on DrivAerNet. Table  1 presents a comprehensive comparison between TripNet and previous models.
@@
These contributions highlight the effectiveness and scalability of our approach for high-fidelity aerodynamic analysis. To the best of the authors’ knowledge, this is the first implementation of the triplane representation to solve partial differential equations (PDEs), as all existing methods have exclusively utilized it for generative modeling.
@@
Figure 1: Illustration of Triplane Networks and their role in CFD tasks. The figure demonstrates how triplane features encode spatial information for aerodynamic analysis, showcasing their applications in predicting surface pressure, wall shear stress, and full 3D flow fields.
Table 1: A comprehensive comparison of various 3D geometric deep learning models for solving Navier-Stokes equations.
Model CFD Tasks Geometry Representation Prediction Prediction Beyond Memory Cost
$C_{d}$ $P$,$\tau_{w}$ $U$ Representation Type Fashion Input Points Independent on Resolution
ResNeXt Song et al. (2023) $\checkmark$ $\times$ $\times$ Images Explicit Images $\times$ $\times$
PointNet Qi et al. (2017a) $\checkmark$ $\checkmark$ $\times$ Point Cloud Explicit Input Nodes $\times$ $\times$
MeshGraphNet Pfaff et al. (2020) $\checkmark$ $\checkmark$ $\times$ Mesh Explicit Graph Nodes $\times$ $\times$
X-MeshGraphNet Nabian (2024) $\checkmark$ $\checkmark$ $\checkmark$ Mesh $+$ Point Cloud Explicit Graph Nodes $\times$ $\times$
Transolver Wu et al. (2024a) $\checkmark$ $\checkmark$ $\checkmark$ Point Cloud $+$ SDF Explicit Input Nodes $\times$ $\times$
FigConvNet Choy et al. (2025) $\checkmark$ $\checkmark$ $\times$ Point Cloud Explicit Input Nodes $\times$ $\times$
GINO Li et al. (2024) $\checkmark$ $\checkmark$ $\checkmark$ Point Cloud $+$ SDF Explicit Query $\checkmark$ $\times$
TripNet (Ours) $\checkmark$ $\checkmark$ $\checkmark$ Triplane Implicit Query $\checkmark$ $\checkmark$
@@ 2.1 Triplane Representation
Triplane representation has recently emerged as a powerful approach in 3D representation learning for scene reconstruction Wu et al. (2024b), 3D geometry Zhang et al. (2024) and avatar generation  Wang et al. (2023); Sun et al. (2023); Zhang et al. (2025) because of its balance between computational efficiency and ability to capture fine-grained details in 3D objects. Three axis-aligned and orthogonal feature planes are used to encode 3D information, offering a compact but expressive way to handle complex 3D shapes and textures. Models such as GANs and diffusion models  Shue et al. (2023); Chan et al. (2022) can be trained on these triplane representations to generate 3D neural fields with high diversity and quality. Building on this foundation, the Roll-out diffusion network (Rodin) Wang et al. (2023) rolls out the triplanes into a single 2D feature plane and performs 3D-aware diffusion to generate 3D digital avatars. Compared to 1D latent vector, voxel, or grid-based representations, triplane representations offer greater memory efficiency while preserving spatial information, providing a 3D-aware encoding that better captures geometric and structural details.
In this work, we aim to convert the 3D car designs to triplanes for TripNet training. Firstly we build the occupancy field for each 3D object by sampling 5 million points uniformly in the domain and another 5 million points near the surface. Next, we followed the two-stage fitting method NFD Shue et al. (2023) used to get the triplanes. Each 3D object is represented by triplanes $\mathbf{P}_{xy},\mathbf{P}_{yz},\mathbf{P}_{xz}\in\mathbb{R}^{C\times H\times W}$. In the fitting, we project the coordinate $\mathbf{x}\in\mathbb{R}^{3}$ onto the triplanes and aggregate the resulting feature vectors to predict the value of occupancy field through an MLP decoder ($MLP_{\Phi}$ with parameters $\Phi$). The occupancy function $occ:R^{3}\rightarrow\{0,1\}$ is defined as:
@@
To mitigate potential compression artifacts in triplane-based 3D representations, we draw inspiration from Neural Field Diffusion (NFD) Shue et al. (2023), where regularization terms are introduced to simplify the data manifold that the diffusion model must learn. It has been shown that these regularization terms effectively reduce significant artifacts in 3D asset generation. We incorporate the same regularization terms in our method to enhance the quality of the triplanes. For further details, interested are encouraged to refer to the original paper Shue et al. (2023). The training objective of the 3D reconstruction task for triplane fitting is as follows:
@@ 2.2 Drag Coefficient Task
To predict the drag coefficient ($C_{d}$), we employ a lightweight convolutional neural network (CNN) to efficiently process triplane representations222model architecture can be found in Figure 3 in the Appendix. The input to our model consists of raw triplane features $\mathbf{P}=[\mathbf{P}_{xy},\mathbf{P}_{yz},\mathbf{P}_{xz}]\in\mathbb{R}^{3\times 32\times 128\times 128}$. Since each feature plane has a channel depth of 32, stacking the three triplanes and taking it as the input would make the model large and inference slow. We apply channel-wise $max$, $min$, and $mean$ pooling on the raw triplane. Each pooling operator $\Theta:\mathbb{R}^{32\times 128\times 128}\rightarrow\mathbb{R}^{3\times 128\times 128}$ compresses the channel depth of original triplane $\mathbf{P}_{ij}$ from 32 to 3. Three feature planes $\mathbf{\tilde{P}}_{ij}$ where $ij\in\{xy,yz,xz\}$ are stacked channel-wise for drag coefficient prediction: